OpenStudy (anonymous):

What is the solution of the equation? 6x – 3 = 5x – 5

OpenStudy (ganpat):

6x - 3 = 5x - 5 bring all x variables on one side and rest on other.. so, 6x - 5x = -5 + 3 x = -2... is the answer

OpenStudy (anonymous):

okie dokie brody

OpenStudy (ganpat):

wow, whats that ???

OpenStudy (anonymous):

haha nvm yo lol

OpenStudy (ganpat):

Lol !! :D

OpenStudy (anonymous):

aye go see if you can answer my other ? lol